Find the integrals of the Pfaffian differential equation
z(z-y)dx+z(x+z)dy+x(x+y)dz=0
1
Expert's answer
2020-12-20T16:50:35-0500

Given differential equation is

"z(z-y)dx +z(x+z)dy+x(x+y)dz = 0"


Let "\vec{X} = z(z-y)\hat{i} +z(x+z)\hat{j}+x(x+y)\hat{k}"

To check whether it is integral ,"\vec{X} .(\nabla\times{\vec{X}}) = 0"

Now, "\nabla\times{\vec{X}} = \begin{vmatrix}
 \hat{i} & \hat{j} &\hat{k}\\
 \frac{\partial }{\partial x} & \frac{\partial }{\partial y} &\frac{\partial }{\partial z}\\
 z(z-y) & z(x+z) & x(x+y)

\end{vmatrix}"



"\nabla\times{\vec{X}} = -2z\hat{i} +2(x+y-z)\hat{j} + (z+1)\hat{k}"

Now, "\vec{X} .(\nabla\times{\vec{X}}) =[ z(z-y)\hat{i} +z(x+z)\hat{j}+x(x+y)\hat{k}
] .[-2z\hat{i} +2(x+y-z)\hat{j} + (z+1)\hat{k}] = 0"

Hence it is integrable.


Let "y = ux \implies dy = xdu + udx"

and "z = vx \implies dz = vdx+xdv"

Then partial differential equation will be reduced to the form

"vx(vx-ux)dx + vx(x+vx)(udx+xdu) + x(x+ux)(vdx+xdv) = 0"


It can be further reduced to

"\frac{dx}{x} + \frac{du}{1}+\frac{dv}{v}-\frac{du}{(u+2v+vu)}-\frac{dv}{(u+2v+uv)} = 0"

Integrating it, we get


"lnx+u+lnv-\frac{ln(2v+u(1+v))}{v+1} -\frac{ln(u+v(2+u))}{2+u} = C"

putting values of u and v,


"lnz+\frac{y}{x}-\frac{ln(2\frac{z}{x}+\frac{y}{x}(1+\frac{z}{x}))}{v+1} -\frac{ln(\frac{y}{x}+\frac{z}{x}(2+u))}{2+\frac{y}{x}} = C= C(x+y+z)"